### v3

# Just Functioning as a Hook for Two-Stage Referring Multi-Object Tracking

Abstract
Referring Multi-Object Tracking (RMOT) aims to localize target trajectories in videos specified by natural language expressions. Despite recent progress, the intrinsic relationship between the two subtasks of tracking and referring in RMOT has not been fully studied. In this paper, we present a systematic analysis of their interdependence, revealing that current two-stage Referring-by-Tracking (RBT) frameworks remain fundamentally limited by insufficient modeling of subtask interactions and inflexible reliance on semantic alignment modules like CLIP. To this end, we propose JustHook, a novel two-stage RBT framework where a Hook module is firstly designed to redefine the linkage between subtasks. The Hook is built centered on grid sampling at the feature-level and is used for context-aware target feature extraction. Moreover, we propose a Parallel Combined Decoder (PCD) that learns in a unified joint feature space rather than relying on pre-defined cross-modal embeddings. Our design not only enhances the interpretability and modularity but also significantly improves the generalization. Extensive experiments on Refer-KITTI, Refer-KITTI-V2, and Refer-Dance demonstrate that JustHook achieves state-of-the-art performance, improving the HOTA by +6.9% on Refer-KITTI-V2 with superior efficiency. Code will be available soon.
Introduction
By integrating natural language processing into traditional Multi-Object Tracking (MOT), Referring Multi-Object Tracking (RMOT) [33] aims to achieve consistent tracking of specific targets through precise guidance provided by natural language. Several recently proposed datasets and studies have revealed the progress. However, a deeply systematic task analysis is still lacking to support the development of a generalizable methodology. In this paper, we delve into a critical yet overlooked issue in RMOT: As a typical cross-task problem, how should the tracking task and referring task be effectively associated?
Figure 1: (a) Referring-by-Tracking classifies MOTR [35]-decoded queries via a linear layer. (b) Tracking-by-Referring associates trajectories within boxes located by GroundingDINO [20].
Although rarely discussed explicitly, as shown in Fig. 1, the underlying logic between subtasks has always been a central focus of research in RMOT. First introduced by TransRMOT [33] and extended by TempRMOT [39], the one-stage Referring-by-Tracking (RBT) paradigms pioneer an end-to-end design that prioritizes the tracking task, where matching scores are computed directly using trajectory queries decoded by MOTR [35]. In contrast, LaMOTer [17], highlights the referring task and constructs trajectories by linking detection boxes predicted by GroundingDINO [20] (i.e. Tracking-by-Referring, TBR). Despite being designed for different scenarios (TempRMOT for closed-set and LaMOTer for open-set), both methods achieve impressive performance within their respective domains.
Figure 2: Comparison of subtask logic in existing methods. (a) previous two-stage RBT reuses the encoder to joint subtask and computes CLIP-based cosine similarity as matching score. (b) Our method directly hooks features from the visual backbone and decodes scores via PCD in a composed feature space.
However, both MOTR [35] and GroundingDINO [17] often suffer from significant computational overhead. Among existing methods, the two-stage RBT paradigm, exemplified by iKUN [7], addresses efficiency by first employing lightweight off-the-shelf trackers to generate trajectory candidates, followed by an independent referring stage as shown in Fig. 2 (a). Nevertheless, this approach has gradually lost its popularity due to its poor generalization performance on complex datasets.
Rethinking these paradigms from the perspective of inter-subtask, we observe that both the popular one-stage RBT paradigm and the TBR paradigm, despite their differing logics, adhere to a common principle: ensuring the sufficiency and necessity of information transmission between preceding and succeeding tasks. Sufficiency implies that the subsequent task can receive all the information it needs, while necessity ensures that the preceding task transmits only the necessary information, without redundancy. In one-stage RBT, the end-to-end design provides a smooth gradient flow, allowing the trajectory queries to adaptively adjust the transmitted information according to downstream demands. In TBR, although lacking continuous gradient propagation, it reduces redundancy by placing the tracking task, which naturally requires less information, after the referring stage. In contrast, current two-stage RBT approaches, despite improving efficiency, only achieve suboptimal sufficiency and low necessity. Unlike the inherent advantages of other paradigms, current two-stage RBT methods can neither offer seamless information flow between subtasks nor eliminate the need to prioritize a lightweight tracker for efficiency. Therefore, during bridging, given that only trajectory boxes and raw images are available as prior information, and meanwhile, the referring subtask involves both local attributes (e.g., color, category) and global spatial cues (e.g., position, direction), these methods have to reuse the encoder to process both full images and cropped target patches via CLIP [28] as illustrated in Fig. 2(a), thereby focusing on the specific target while enriching the spatial context.
However, a widely known fact but neglected by them is that most visual backbones (e.g., ResNet [12], Swin [22]) inherently possess strong spatial contextual modeling capabilities. Even CLIP, though pre-trained at the image-level, has demonstrated fine-grained recognition abilities [31; 16]. Hence, neglecting the pre-existing capabilities and instead introducing additional re-training modules is a suboptimal approach to achieving sufficient information propagation. The irrelevant content in global images and the linking of subtasks by reusing the encoder also undermine the necessity. The 38.52% drop in HOTA of iKUN [7] on the Refer-KITTI-v2 [39] benchmark further demonstrates that unnatural information flow fails to fully exploit the backbone’s potential and compromises its robustness.
Therefore, rethinking the coupling between subtasks in a more deliberate and holistic perspective, a better bridging process should (1) precisely extract target regions by using tracking boxes to maximize necessity, and (2) preserve backbone’s spatial gradient flow to obtain contextual information and achieve better sufficiency. Based on these two principles, we find that nothing fits them better than the grid sampling operation. Therefore, centered around it, we build a new two-stage RBT framework, named JustHook, whose name draws inspiration from hook functions in programming [23]. Although grid sampling has been widely used in other methods such as ROI pooling [29] and Deformable Attention [43], its role in RMOT remains unexplored. To the best of our knowledge, we are the first to identify its suitability for RMOT and unlock its potential through carefully designed supporting modules.
Furthermore, to leverage the advantages of the new bridging module on dense visual backbones, we propose an active combined feature metric to replace the traditional cosine similarity matching by introducing the Parallel Combined Decoder (PCD). Finally, JustHook can be coupled with any backbone and text encoder and achieve robust performance. Our contributions can be summarized as follows:
(1) We offer a new perspective—subtask correlations, to reevaluate existing RMOT paradigms and reveal the inherent flaws in current two-stage RBT frameworks.
(2) Rethinking the subtask correlation patterns in the two-stage RBT paradigm, we propose a new framework, JustHook, which establishes well-suited information pathways and enables flexible coupling with arbitrary encoders.
(3) JustHook outperforms SOTA methods on the Refer-Dance [7], Refer-KITTI [33] and Refer-KITTI-V2 [39] datasets. It achieves a notable 6.9% HOTA improvement on Refer-KITTI-V2 and performs the highest efficiency.
Related Work
Figure 3: The overall framework of JustHook. During the encoding process of the visual backbone, JustHook directly extracts the target regions from the multi-scale visual features through grid sampling. The optical flow information constructed from these features is then aggregated into the trajectory features $J$ after OME. Finally, the trajectory features and multiple linguistic features are decoded in parallel into multiple groups of Match Score $O$ in PCD. The purple dotted line represents the backpropagation path of the gradient. Even though only the features of the target regions are sampled, the VFH will still propagate the gradient to other regions through visual backbone, thus adaptively obtaining the required contextual information.
Referring Single Object Tracking
Referring Single Object Tracking (RSOT) has been studied for several years. Most SOTA methods [2; 34; 10; 9] follow classic tracking strategies, integrating advancements in referring video segmentation [6] (or detection [32]) and focusing on end-to-end joint tracking algorithms. JointVLT [40] performs grounding in the first frame by inputting only a zero-pad, and then continuously tracks the target in subsequent frames based on template and expression. TransVLT [42] uses a set of proxy tokens for cross-modal information fusion, which project linguistic features and generate kernel vectors to modulate visual features. Instead of directly decoding the bounding boxes, MMTrack [41] constructs a set of queries from expression and bounding box cues and generates the target trajectories using an auto-regressive approach. Although significant achievements [15; 36] made, the development of RSOT is limited by highly limited task scenarios.
Referring Multi Object Tracking
TransRMOT [33] pioneered RMOT by introducing the Refer-KITTI dataset along with a baseline model based on MOTR [35]. Building upon TransRMOT, MLS-Track [26] further enhances cross-modal interaction through a multi-level module design. To address the high computational cost of these methods, iKUN [7] innovatively proposed a modular two-stage RBT paradigm. Recently, TempRMOT [39] constructs the enhanced Refer-KITTI-V2 dataset by leveraging large language models [3]. Focusing on RMOT performance under open-vocabulary scenarios, LaMOTer introduces a Tracking-by-Referring architecture to further enrich the subtask logic of RMOT. While each paradigm has its own strengths and weaknesses, few studies have examined the internal interaction between subtasks in RMOT or identified the limitations of current two-stage RBT methods.
Method
Problem Formulation
In RMOT, the trajectory of target $i$ can be represented by its sequence of detection boxes across frames, denoted $\mathcal{B}^{i}=\{B^{i}_{t}\}_{t=t^{i}}^{T^{i}}$, where $t$ indicates the frame index, and $t^{i}$ and $T^{i}$ denote the starting and ending frames in which the target appears. We use $\mathcal{B}^{i}_{x:y}$ to denote a segment of target $i$’s trajectory, where $x$ and $y$ represent the starting and ending frames of this segment, respectively. The index $t$ is not necessarily continuous, as the target may disappear or be occluded in certain frames and reappear later. For each video, let $\mathcal{O}=\{\mathcal{B}_{i}\}_{i=1}^{K}$ be the set of all $K$ target trajectories, and let $\mathcal{E}=\{E_{j}\}_{j=1}^{N}$ denote the set of $N$ natural language expressions annotated for the targets in the whole video. The process of inferring all target trajectories $\mathcal{O}$ from the video frame sequence $\mathcal{I}=\{I_{t}\}_{t=0}^{T}$ constitutes the tracking subtask, while the referring subtask can be formulated as learning a many-to-many matching relation between the expression set $\mathcal{E}$ and the trajectory set $\mathcal{O}$:
$\mathcal{R}\subseteq\mathcal{E}\times\mathcal{O}$
where $(E_{j},\mathcal{B}^{i}_{x:y})\in\mathcal{R}$ indicates that expression $E_{j}$ refers to trajectory segment $\mathcal{B}^{i}_{x:y}$. Noticed that the reason why we use a segment instead of the entire trajectory is that certain expressions, such as “people on the left”, may become invalid as the target moves. In the two-stage RBT paradigm, the tracking subtask is handled by an off-the-shelf tracker, and we focus solely on the bridging module and the referring subtask.
Overview
To facilitate understanding, as illustrated in Fig. 3, we decompose JustHook into three main components: (1) Hook, a coupling module centered around grid sampling; (2) OME, an Optical-flow-based Motion Extractor; and (3) PCD, a Parallel Combined Decoder. Given a $p$-frame segment of a trajectory, denoted $\hat{\mathcal{B}}_{t:t+p}$, we sample $\hat{N}$ expressions $\hat{\mathcal{E}}=\{E_{j}\}_{j=1}^{\hat{N}}$ from the expression set $\mathcal{E}$ during training, while during inference, we iterate over all expressions. JustHook ultimately outputs $\hat{N}$ matching scores per forward, each corresponding to the match between the trajectory segment $\hat{\mathcal{B}}_{t:t+p}$ and one of the expressions in $\hat{\mathcal{E}}$.
Hook
Serving as a bridging module, Hook is designed to receive the output from the tracking subtask, denoted as $\hat{\mathcal{B}}_{t:t+p}$, and extract features required by the referring subtask. Unlike previous approaches that rely on additional modules or encoding procedures, Hook directly grid-samples the target features from the raw encoding stream of the backbone at the feature level. Its smooth gradient path makes full use of the contextual aggregation capability of pretrained models. Hook consists of two components: a neighboring grid sampling specifically adapted for RMOT, and a long-range modeling enhancement module.
Neighboring Grid Sample.
Naive grid sampling typically relies on hand-crafted or learnable coordinate grids. In RMOT, however, the coordinate grid can be naturally obtained from the tracking results $\hat{\mathcal{B}}_{t:t+p}$. Given a coordinate grid $P$, the value at any point $X$ is obtained by interpolating its four surrounding feature values, denoted $Y_{a},Y_{b},Y_{c},$ and $Y_{d}$. The interpolation weights $D_{a},D_{b},D_{c},$ and $D_{d}$ depend on the offsets, and satisfy $D_{a}+D_{b}+D_{c}+D_{d}=1$. Taking $Y_{a}$, the top-left neighbor of $X$, as an example: if its normalized horizontal and vertical offsets relative to $X$ are $(u,v)$, then its interpolation weight is $D_{a}=(1-u)(1-v)$. $X$ can be calculated as follows:
$X=D_{a}\cdot Y_{a}+D_{b}\cdot Y_{b}+D_{c}\cdot Y_{c}+D_{d}\cdot Y_{d}$
While naive grid sampling is already well-aligned with the requirements of the bridging module, a notable distributional gap exists in the two-stage RBT paradigm: the referring subtask is typically trained with ground-truth trajectories, whereas inference relies on trajectory candidates produced by an off-the-shelf tracker. To narrow this gap, we perturb the coordinate grids to simulate potential tracking errors. Specifically, we adopt three operations (i) randomly sample fragmented trajectory segments to emulate target loss, (ii) inject Gaussian noise to model localization inaccuracies, and (iii) re-combine grid sequences within a batch to mimic identity switches. This augmented strategy enables the model to acquire neighborhood awareness across temporal, spatial, and instance dimensions, thereby reducing overfitting. Furthermore, spatial neighbor sampling enhances the model’s capacity to capture broader contextual information.
Figure 4: Illustration of Long-Range Enhancement. Guided by the linguistic feature $F_{l}$, we computes $M$ reference points $P_{r}$ through a Transformer decoder and a residual fully connected network, which then are used to actively provide reference features by grid sampling.
Long-Range Enhancement.
Although grid sampling ensures a smooth gradient path, the model’s perception of external features will weaken as the spatial distance from the sampled region increases. To address this, we design an additional long-range modeling enhancement module for Hook to help the model actively acquire global features.
As illustrated in Fig. 4, we incorporate linguistic priors by leveraging expression information and perform cross-attention between learnable query vectors $Q_{LR}\in\mathbb{R}^{\hat{N}\times M\times C}$ and the linguistic features $F_{l}\in\mathbb{R}^{\hat{N}\times L\times C}$ output by the text encoder. The resulting attention output is then passed through an multi-Layer perceptron (MLP) and a sigmoid layer to generate normalized 2D reference points $P_{r}$. These reference points are used alongside the coordinate grid $P$ to extract visual features via grid sampling. The sampled features are then temporally pooled to form the reference feature $F_{r}\in\mathbb{R}^{\hat{N}\times M\times C}$. Finally, in PCD, the reference features are fused with the combined features to compute the matching score.
Optical-flow-based Motion Extractor
Motion information is an indispensable part of processing motion expressions (e.g., “the car turning left”) in RMOT. However, the existing methods only use simple temporal pooling to compress multi-frame information [7], ignoring the target optical flow naturally provided by the trajectory proposals. Therefore, we concatenate along the channel dimension both the coordinate grid displacements $\Delta Grid=Cat(\{P_{t-i}-P_{t-i-1}\}_{i=0}^{T})$, which serve as explicit optical flow (the blue part in Fig. 5), and the multi-frame features $F_{v}\in\mathbb{R}^{p\times C\times H\times W}$, which serve as feature optical flow (the purple part in Fig. 5). To further enhance the model’s perception capability in the spatial dimension, we design three different strategies: (1) MLP strategy, where features are simply fused using a multi-Layer perceptron (MLP); (2) Conv strategy, which applies convolution after channel-wise concatenation to focus on the local information of the target; and (3) Attention strategy, which leverages inter-frame self-attention [30] to adaptively highlight regions of interest.
Figure 5: Three strategy of OME. (a) MLP strategy only transforms channel, (b) Conv strategy perceives space using convolution, and (c) Attention strategy perceives space using self-attention.
Parallel Combined Decoder
The current two-stage RBT paradigm usually takes CLIP [28] as a mandatory component, using its inherent cross-modal alignment capability to directly measure the matching between trajectory features $F_{j}$ and linguistic features $F_{l}$ via cosine similarity. On one hand, the theoretical upper bound of CLIP limits the potential of previous methods. On the other hand, our grid-sampling-based hook tends to perform better with backbones that excel in dense and long-range modeling scenarios. Therefore, as illustrated in Fig. 6 (b), we introduce a Parallel Combined Decoder (PCD) that operates on a joint embedding space formed by concatenating visual features and expression words, allowing the model to leverage the inherent differences between positive and negative pairs instead of relying on hard-to-measure cross-modal similarity in a misaligned space.
Specifically, as shown in Fig. 6, PCD concatenates the trajectory features $F_{j}\in\mathbb{R}^{C\times H\times W}$ and linguistic features $F_{l}\in\mathbb{R}^{\hat{N}\times L\times C}$ in the spatial dimension (the word dimension in $F_{l}$), and then uses them as the Key and Value, with the learnable query vectors $Q\in\mathbb{R}^{\hat{N}\times C}$ as the Query, to perform masked cross-attention. Under the long-range enhancement setting, we further include the reference features $F_{r}\in\mathbb{R}^{\hat{N}\times M\times C}$ in the concatenation to actively extract global information. The decoded queries are then further summarized through a Feed-Forward Network (FFN) followed by an MLP to generate the match scores $O\in\mathbb{R}^{\hat{N}\times 2}$.We control the cross-attention mask $A\in\mathbb{R}^{\hat{N}\times(H\cdot W+\hat{N}\cdot L+\hat{N}\cdot M)}$ so that each query vector can attend only to its corresponding textual features and reference features, as well as all visual features.
Table 1: Comparison with state-of-the-art methods on Refer-KITTI [33] and Refer-KITTI-V2 [39]. “” indicates that long-range enhancement is not used.
Method HOTA DetA AssA LocA
Refer-KITTI
One-Stage DeepRMOT [13] 39.55 30.12 53.23 80.49
TransRMOT [33] 46.56 37.97 57.33 90.33
MLS-Track [26] 49.05 40.03 60.25 -
TempRMOT [39] 52.21 40.95 66.75 90.40
Two-Stage ByteTrack [37] 24.95 15.50 43.11 73.90
CStrack [18] 27.91 20.65 39.10 79.51
TransTrack [33] 32.77 23.31 45.71 79.48
TrackFormer [27] 33.26 25.44 45.87 79.63
iKUN [7] 48.84 35.74 66.80 -
JustHook CLIP [28] CLIP-R 50 [28] 53.49 43.62 65.68 90.90
RoBERTa [21] Resnet34 [12] 53.47 43.01 66.55 90.95
ROPE Swin-T [14] 53.91 44.02 66.09 90.97
Refer-KITTI-V2
One-Stage TransRMOT [33] 31.00 19.40 49.68 89.82
TempRMOT [39] 35.04 22.97 53.58 90.07
Two-Stage FairMOT [38] 22.53 15.80 32.82 78.28
ByteTrack [37] 24.59 16.78 36.63 78.00
iKUN [7] 10.32 2.17 49.77 74.56
JustHook CLIP [28] CLIP-R 50 [28] 41.03 29.08 58.02 89.03
RoBERTa [21] Resnet34 [12] 41.50 29.21 59.10 89.25
ROPE Swin-T [14] 41.94 30.45 57.92 89.00
Figure 6: Illustration of the PCD. Compared with the passive strategy of cosine similarity, PCD actively examines the matching of combined sample pairs by concatenation and conducts contrastive learning, no longer relying on the modality alignment capability inherent in CLIP.
Loss Computation
We apply deep supervision to JustHook and supervise all intermediate and final outputs using Focal Loss [19], to enhance the model’s multi-scale capability and mitigate the imbalanced sample distribution in the dataset.
Experiment
Experiment Setup
Dataset and Metrics.
We report the performance of JustHook on 3 datasets: Refer-Dance [7], Refer-KITTI [33] and Refer-KITTI-V2 [39]. For a fair comparison, we use Higher Order Tracking Accuracy [25] (HOTA) as the main metric, following TransRMOT [33]. HOTA comprehensively considers the detection and association capabilities of the tracking model. We also report the localization accuracy (LocA), the detection accuracy (DetA) and the association accuracy (AssA).
Implement Details.
As a two-stage RBT method, JustHook allows flexible replacement of any detector/tracker without the need for retraining. Therefore, we report only the results using the best-performing detector and tracker: TempRMOT [39] and NeuralSORT [7]. TempRMOT [39], used as the detector, is trained with detection annotations by treating each object in each frame as a new target. Structurally, it can be viewed as a Deformable-DETR [43] augmented with inter-frame interactions. NeuralSORT [7], built upon the classical SORT [1] framework, replaces the Kalman filter’s observation and transition models with neural networks, achieving excellent tracking performance. The impacts of different detectors and trackers [4; 8; 7; 43] on JustHook can be found in Supplementary Materials.
In addition, JustHook can also connect with any visual backbone and text encoder. The visual backbones we used include CLIP [28], ResNet [12] and ROPE Swin [14], all of which use the pre-trained parameters on ImageNet [5], with an input size of $224\times 672$ to align as closely as possible with the aspect ratio of video. For text encoders, we test both CLIP [28] and RoBERTa [21].
During training, we iterate the trajectory segment with sample stride 2, and each sample consists of 4 frames. In neighborhood grid sampling, to simulate target loss, we sample 8 frames each time and randomly select 4 of them. For localization noise, we apply Gaussian noise with zero mean and a standard deviation equal to 0.02% of the original size to the width, height, and center of each bounding box. Random frame re-combination is performed with a probability of 1%. During inference, the sample stride is set to 1. $\hat{N}$ and $M$ are set to 36 and 10, respectively. We use AdamW [24] as the optimizer, with a learning rate of 3e-5, and train for 20 epochs. All experiments are performed on 2 NVIDIA 4090 GPUs.
Benchmark Experiments
Table 2: Performance comparison on Refer-Dance [7].
Method HOTA DetA AssA
TransRMOT [33] 9.58 4.37 20.99
iKUN [7] 29.06 25.33 33.35
JustHook-best 31.83 30.31 33.47
We compare our method with SOTA approaches on Refer-KITTI and Refer-KITTI-V2 in Table 1. As discussed in the following sections, we do not apply additional long-range enhancement to the Transformer-based backbone. JustHook achieves the best performance under various encoder configurations. Compared to the current SOTA method TempRMOT [39], our best model, JustHook-best, which adopts RoBERTa [21] as the language encoder and ROPE Swin-T [14] as the visual backbone, achieves superior performance on both Refer-KITTI and Refer-KITTI-V2. JustHook demonstrates a more significant advantage on Refer-KITTI-V2, surpassing TempRMOT by 6.9%, 7.48%, and 4.34% on HOTA, DetA, and AssA, respectively. We further report the performance of JustHook-best on Refer-Dance in Table 2, which indicates that JustHook still exhibits robust performance even in non-autonomous driving scenarios.
Table 3: Comparison of training and inference time. All experiments are conducted on Refer-KITTI-V2 [39]. “”: iKUN [7] trains at the trajectory level. The detector used is the modified TempRMOT.
Method Epoch Train Detect Infer FPS HOTA
TempRMOT [39] 50 62 h - 299 min 25.5 35.04
iKUN [7] 100 2.51 h 2.72 min 115 min 64.7 10.32
JustHook-best 20 2.17 h 2.72 min 69 min 106.3 41.94
Qualitative Results.
The qualitative results can be found in the Supplementary Materials.
Ablation Study
Table 4: Comparison of different spatial perception strategies in OME.
Strategy HOTA DetA AssA
Avg-Pool 40.69 28.14 58.97
Attention 41.36 28.96 59.20
Conv 41.51 29.42 58.71
MLP 41.94 30.45 57.92
OME.
We compare the performance of OME under Just-best setting, where Avg-Pool refers to the temporal average pooling used in iKUN [7] as a replacement for OME. As shown in Table 4, OME extracts motion information more effectively than naive pooling. However, it seems that more complex aggregation methods perform even worse than the simple MLP-based approach. This can be attributed to the fact that pre-trained visual backbones already encode rich static information within each frame. Overly complex aggregation strategies may distort the feature space, thereby hindering subsequent learning.
Table 5: Comparison of the performance about Hook and PCD on Refer-KITTI-V2.
Method Hook PCD HOTA DetA AssA
iKUN 10.32 2.17 49.77
JustHook-CLIP 35.88 22.39 57.61
JustHook-CLIP 40.14 28.26 57.16
Hook and PCD.
In Table 5, we evaluate the performance of Hook and PCD on Refer-KITTI-V2. Here, we use CLIP as encoder, and both Long-Range Enhancement and OME are disabled for all models for a fair comparison with iKUN. ✘ indicates that cosine similarity is used directly to compute the matching score instead of PCD. The results show that both PCD and Hook bring significant performance improvements. Moreover, combining with the performance of JustHook in different encoder settings, PCD can be found to perform well not only in aligned embedding spaces but also maintain strong performance in non-aligned spaces, demonstrating its potential to replace cosine similarity.
Neighboring and Long-Range Enhancement.
Table 6: Comparison of the performance about Neighboring and Long-Range Enhancement
Method $M$ HOTA DetA AssA
JustHook-CLIP 0 40.59 27.87 59.23
10 41.03 29.08 58.02
30 40.85 28.91 57.87
JustHook-R 34 0 40.97 29.85 56.39
10 41.50 29.21 59.10
30 41.39 29.14 58.92
JustHook-best 0 41.94 30.45 57.92
10 41.63 29.95 58.01
30 41.56 30.00 57.71
JustHook-best
w/o Neighbor
0 40.65 28.86 57.42
We analyze the impact of Neighboring Grid Sample and Long-Range Enhancement on Refer-KITTI-v2 in Table 6, where $M$ denotes the number of sampling points used for Long-Range Enhancement, with 0 indicating no enhancement. As observed, convolution-based visual backbones exhibit better adaptability to long-range enhancement compared to Transformer-based ones. This is because of the inherent strength of Transformers in modeling long-range dependencies, thus weakens the effect of our additional active enhancement. The number of reference points has a relatively minor impact on model performance, and setting $M=10$ is an empirical choice. In the setting without enhancement, removing the noise introduced by Hook leads to a performance drop, highlighting the importance of simulating the noise from the tracking task.
Training and inference time.
Table 3 shows a comparison of efficiency on the same hardware. Both iKUN and TempRMOT follow the official configuration. Since iKUN only samples 4 frames of each trajectory for training in each epoch, its total epochs are significantly high. When training the detector, as the dataset contains only two “expressions” (car and pedestrian), our detector TempRMOT can converge in fewer than 10 epochs, instead of the 50 epochs reported in its RMOT version. The speed is generally faster in other detectors such as Deformable-DETR [43]. Considering that most TBD trackers are training-free, even when accounting for the time spent training the first stage, JustHook still demonstrates a significant efficiency advantage.
Conclusion
In this work, we address the limitations of existing two-stage RBT methods by proposing JustHook, a flexible and efficient hook-like framework. JustHook introduces two novel components, Hook and PCD, which enable context-aware feature aggregation and parallel decoding in a combined space, achieving significant gains in accuracy and computational efficiency. By functioning like a hook, JustHook efficiently leverages the inherent contextual aggregation capabilities of pre-trained models and introduces a novel encoder-decoder architecture for the two-stage RMOT, demonstrating strong adaptability to various emerging visual/text encoders and ensuring future-proofing.
Dataset Details
Refer-KITTI
Refer-KITTI [34] uses 18 low-complexity videos and tracking annotations from the KITTI [11] training set, encompassing a total of 895 expressions [39]. During the annotation process, manual bidirectional verification is performed on the matching between expressions and targets to ensure annotation accuracy. Although Refer-KITTI lays the foundation for RMOT, its test set is relatively small and exhibits a significant distribution gap from the training set, resulting in a failure to accurately reflect real-world scenarios [39].
Refer-KITTI-V2
Recognizing the issues in Refer-KITTI, Refer-KITTI-V2 [39] leverages large language models to further enrich the diversity of expressions in Refer-KITTI. Building on the 18 videos from Refer-KITTI, Refer-KITTI-V2 includes 9758 expressions and an additional 3 videos from the KITTI [11] training set, with 2 used for training and 1 for testing. The diversified expressions and more complex scenarios make Refer-KITTI-V2 more reflective of real-world conditions compared to Refer-KITTI, presenting new challenges for RMOT.
Correction for Refer-KITTI
Refer-KITTI has suffered from multiple annotation errors for a long time. Many critical issues such as frame misalignment have been addressed in iKUN [7] and Refer-KITTI-V2 [39]. However, the problem of incomplete reference persists in the current version. For instance, in frames 231-232 of video “0002”, two silver vehicles are referenced under the label “left cars in light color” but are missing in the “left cars in silver”. In addition to these general errors, there are also some special errors in Refer-KITTI, and the impacts they have on one-stage methods and two-stage methods are inherently inconsistent. That is, numerous objects originally annotated in the tracking annotations of Refer-KITTI are entirely omitted in referring annotations of Refer-KITTI, such as all pedestrians in video “0001”. Because one-stage methods use all tracking annotations for training while two-stage methods are only based on referring annotations, this issue leads to particularly significant adverse effects on two-stage methods.
Table 7: Performance comparison of TempRMOT [39] before and after the correction of Refer-KITTI [33].
Correction HOTA DetA AssA
Before 52.21 40.95 66.75
After 51.79 40.28 66.68
Table 8: Performance comparison of multi-object tracking on KITTI [11] (using the same split as Refer-KITTI-V2 [39]). “” and “††” denote the tracker and detector versions of TempRMOT [39], respectively.
Detector Tracker HOTA DetA AssA LocA
One-stage TempRMOT 63.47 56.09 72.04 91.19
Two-stage D-DETR [43] OC-SORT [4] 54.13 50.31 58.46 90.73
StrongSORT [8] 56.98 54.13 60.20 90.24
NeuralSORT [7] 56.84 54.37 59.67 89.89
TempRMOT†† OC-SORT [4] 58.26 55.35 61.57 90.44
StrongSORT [8] 59.42 57.25 61.92 89.91
NeuralSORT [7] 61.10 57.56 65.08 89.84
Therefore, for a fair comparison, we select expressions from Refer-KITTI-V2 that consist solely of words previously used in Refer-KITTI and replace the original annotations with them. Due to the complete referencing of these missing targets in Refer-KITTI-V2, this approach resolves the existing issues in Refer-KITTI without introducing new words or causing data leakage. To verify the rationality of our correction strategy, we retrain TempRMOT [39] using the corrected Refer-KITTI dataset. As shown in Table 7, our correction strategy does not affect the one-stage methods, since TempRMOT itself already utilizes all the tracking annotations.
Table 9: Performance comparison of JustHook coupled with different trackers, visual backbones, and text encoders on Refer-KITTI-V2 [39].
Detector
& Tracker
Text
Encoder
Visual
Backbone
HOTA DetA AssA LocA
TempRMOT [39] 35.04 22.97 53.58 90.07
D-DETR [43]
& OC-SORT [4]
RoBERTa [21] Resnet34 [12] 38.20 27.95 52.34 89.91
ROPE Swin-T [14] 37.94 28.59 50.53 89.72
CLIP [28] CLIP [28] 36.13 26.31 49.75 89.99
D-DETR [43]
& StrongSORT [8]
RoBERTa [21] Resnet34 [12] 39.70 29.01 54.49 89.53
ROPE Swin-T [14] 39.66 29.85 52.85 89.35
CLIP [28] CLIP [28] 37.44 27.33 51.45 89.65
D-DETR [43]
& NeuralSORT [7]
RoBERTa [21] Resnet34 [12] 39.62 28.86 54.55 89.18
ROPE Swin-T [14] 39.46 29.67 52.64 88.98
CLIP [28] CLIP [28] 37.76 27.23 52.49 89.34
TempRMOT [39]
& OC-SORT [4]
RoBERTa [21] Resnet34 [12] 39.57 29.01 54.13 89.71
ROPE Swin-T [14] 39.69 30.10 52.51 89.47
CLIP [28] CLIP [28] 37.47 27.36 51.46 89.85
TempRMOT [39]
& StrongSORT [8]
RoBERTa [21] Resnet34 [12] 41.03 29.29 57.61 89.33
ROPE Swin-T [14] 41.37 30.51 56.26 89.09
CLIP [28] CLIP [28] 38.89 27.90 54.34 89.49
TempRMOT [39]
& NeuralSORT [7]
RoBERTa [21] Resnet34 [12] 41.50 29.21 59.10 89.25
ROPE Swin-T [14] 41.94 30.45 57.92 89.00
CLIP [28] CLIP [28] 41.03 29.08 58.2 89.03
Discussion and Performance of other Trackers
In order to better illustrate that the gains generated by JustHook do not heavily rely on the improvement of the tracker’s performance, in Table 8, we compare the pure multi-object tracking performance of different detector/tracker combinations (for TBD paradigm) employed in our approach with that of TempRMOT (as a pure one-stage MOT model). Here, “D-DETR” refers to DeformableDETR [43]. In the two-stage (TBD) method, TempRMOT†† serves solely as a detector and adheres to the detection paradigm, where we treat each target in each frame as a new addition. Combined with the comparison of RMOT capabilities in paper, it can be seen that even trackers with weaker tracking capabilities than TempRMOT can achieve performance exceeding the SOTA method TempRMOT [39] in RMOT after integrating JustHook. This demonstrates that JustHook significantly surpasses TempRMOT in matching capability, providing excellent results even when coupled with inferior trackers.
Furthermore, we compare the performance changes of JustHook on RMOT after replacing these TBD-style trackers in Table 9. The version of JustHook is the same as that in the paper. JustHook achieves consistent performance improvements as the capability of the detectors/trackers increases, highlighting its flexibility and robustness.
More experiments
Table 10: Comparison of Performance between iKUN [7] and JustHook when coupled with different encoders on Refer-KITTI [33].
Visual
Backbone
Text
Encoder
HOTA DetA AssA
iKUN [7] CLIP [28] CLIP [28] 48.84 35.74 66.80
ROPE Swin [14] RoBERTa [21] 29.79 14.73 60.38
JustHook CLIP [28] CLIP [28] 53.49 43.62 65.68
ROPE Swin [14] RoBERTa [21] 53.91 44.02 66.09
Impact of Freezing
In Table 11, we evaluate the impact of freezing different components in JustHook, where JustHook-CLIP represents the CLIP-R 50 [28] encoder pair, and JustHook-best refers to the ROPE Swin-T [14] + RoBERTa [21] encoder pair.
It can be observed that different encoders respond differently to parameter freezing. For the CLIP pre-trained encoder, which is naturally aligned across modalities, training only one side disrupts the original embedding space, leading to performance even worse than freezing both sides entirely. In contrast, for RoBERTa and ROPE Swin-T, freezing the text encoder results in a significantly larger performance drop compared to freezing the visual encoder. This can be attributed to the fact that the textual distribution in the Refer-KITTI-V2 dataset differs more from RoBERTa’s pre-training corpus, being focused solely on the autonomous driving, than the visual gap does. Therefore, the model requires more learning on the text side to adapt to this domain shift.
Moreover, Table 11 also shows that JustHook generally outperforms CLIP when using denser visual backbones. However, the efficiency gain brought by freezing the CLIP encoder is unmatched by any of the other encoders. Therefore, by flexibly selecting different encoders based on the actual hardware environment, JustHook demonstrates strong potential for broad real-world applicability.
Table 11: Performance of JustHook after freezing different modules on Refer-KITTI-V2 [39].
Method
Visual
Backbone
Text
Encoder
HOTA DetA AssA
JustHook
-CLIP
Trainable Frozen 40.23 28.96 56.05
Frozen Trainable 40.08 28.44 56.60
Frozen Frozen 40.70 29.03 57.22
Trainable Trainable 41.03 29.08 58.02
JustHook
-best
Trainable Frozen 40.79 28.34 58.88
Frozen Trainable 41.59 29.78 58.20
Frozen Frozen 40.62 27.81 59.46
Trainable Trainable 41.94 30.45 57.92
The Scalability of JustHook
In Table 12, we evaluate the performance of JustHook with different sizes of ResNet [12] and ROPE Swin Transformer-T [14] as visual backbones. The text encoder maintains RoBERTa [21]. It can be observed that as the size of the Swin Transformer increases, the performance of JustHook does not improve significantly and even decreases. However, for convolutional networks, as the size of ResNet increases, the performance of JustHook also improves, but this improvement stops at ResNet34. This indicates that the scalability of visual backbones of different sizes for JustHook is not significant, but CNN-based backbones slightly outperform Transformer-based ones. This can be attributed to the limited visual information of most targets in the current Refer-KITTI-V2 dataset, which does not require a very powerful visual backbone to characterize the target’s state. When the visual backbone is too powerful, the model is more prone to overfitting. Therefore, we finally selected ResNet34 and ROPE Swin-T for reporting.
Table 12: Comparison of JustHook performance with different-sized visual encoders on Refer-KITTI-V2 [39].
Backbone Size HOTA DetA AssA
ROPE Swin [14] Tiny 41.94 30.45 57.92
Small 41.27 30.02 56.89
Base 41.20 28.97 58.72
ResNet [12] 18 40.90 28.80 58.20
34 41.50 29.21 59.10
50 41.17 29.70 57.20
The reliance on CLIP of iKUN and JustHook
To support our viewpoint that iKUN has a strong dependence on the semantic space of CLIP, we quantitatively measure the performance changes of iKUN after replacing the semantic space. Following the official configuration of iKUN [7], we evaluate its performance when coupled with other encoders in Table 10 to explore the adaptability of the current two-stage methods. Here, ROPE Swin [14] uses the tiny version, CLIP [28] adopts the ResNet50 version following iKUN [7]. It is evident that when using the best encoder combination in JustHook (i.e., ROPE Swin-T [14] & RoBERTa [21]), iKUN [7] fails to fully leverage the capabilities of advanced encoders, as JustHook does. When in a semantic space with misaligned modalities, the adaptability of iKUN’s modality-independent passive measurement is significantly lower than the combined space decoding in JustHook, resulting in a significant performance drop and making it difficult to converge. This indicates that iKUN heavily depends on encoder pairs with inherent feature alignment capabilities, such as CLIP [28], and struggles to integrate with emerging encoders, making it not capable of future proofing.
Figure 7: Visualization of some qualitative results (Part 1).
Limitations
Many experiments show that JustHook has a significantly greater gain on Refer-KITTI-V2 than on Refer-KITTI. On the one hand, it can be attributed to the fact that the Refer-KITTI dataset is simpler, with less data and relatively simple text, leaving limited room for improvement. This can be verified in Table 8. Without considering the expressions at all, the pure multi-object tracking performance on Refer-KITTI is only 63.7% HOTA, while the HOTA of over 50 on RMOT nowadays is already very close to this theoretical upper limit. On the other hand, as a two-stage method, JustHook is fundamentally based on solving the classification hyperplane in the combined feature space. This means that compared with one-stage methods, two-stage methods generally have higher requirements for the amount of data and its diversity. Larger quantities of more complex training data are often helpful in providing more support vectors to assist in determining the classification hyperplane. Therefore, JustHook shows a more obvious improvement in performance on the complex dataset Refer-KITTI-V2 and is also more suitable for practical applications in complex real-world scenarios. In the future, through methods such as data augmentation and synthesis, exploring the work of mining difficult samples will be more conducive to the development of two-stage RMOT.
Broader Impact
JustHook further accelerates the research of RMOT, which enriches the technical means of computer vision by using language expressions as semantic cues to guide multi-object tracking predictions. This is of great significance for the development of related industries such as autonomous driving, which can improve the environmental perception ability of autonomous vehicles and enhance the safety and reliability of driving.
Qualitative Results
There are several typical referring results in Fig. 7. In the first expression, JustHook effectively distinguishes between light-colored and dark-colored vehicles and accurately understands the target’s location. In the second expression, despite the continuous movement of the visual capture, JustHook still accurately identifies that the vehicles on the right are stationary. The third expression is more complex, involving both the target’s location and gender. The final expression highlights JustHook’s adjustment in matching the target during movement. In the video, a group of pedestrians cross the road, but the expression only describes pedestrians on the right side. As soon as the target reaches the center or shifts slightly to the left, JustHook cancels the match.
We visualize additional qualitative results in Fig. 8 and Fig. 9 to provide an intuitive comparison between JustHook and TempRMOT [39], where JustHook employs the encoder combination of ROPE Swin-T [14] + RoBERTa [21]. Here, (a) represents the results of TempRMOT [39], (b) represents the results of JustHook, GT denotes the ground truth, and $T_{1}$, $T_{2}$, and $T_{3}$ correspond to 3 sequentially sampled frames (which are not consecutive). In various scenarios, JustHook has demonstrated robust advantages, showing fewer error cases and lost cases compared to TempRMOT [39].
Figure 8: Visualization of some qualitative results (Part 2).
Figure 9: Visualization of some qualitative results (Part 3).